Science Night Interrupt Guidelines

The criteria for interrupting an observing program for a hand-over to the LUCI commissioning team are: Seeing, as measured by the DIMM, reaches 1.1 arc seconds, and for 20 minutes never goes above 1.2. As below, the OSA will contact the LUCI commissioning team at this time. The block will end if the AO system loses lock repeatedly (e.g., due to frame skips, bayside stage timeouts, or rips), the seeing goes to and remains at or above 1.2 arc seconds continuously for 20 minutes, the maximum test time of 4 hours is reached, or the end of the night is reached, whichever comes first. Only one interrupt per night will be requested.

Interruptible Nights:
  • 06 - 17 June (D + OSURC)
  • 26 - 30 June (D + INAF)
  • 01 - 09 July (D + INAF + AZ)

...only the LBTI nights are excluded between now and the summer shutdown. TOO programs are also excluded.

Once the seeing reaches 1.1 arc seconds, and the clock starts ticking on the 20 minutes at less than 1.2 arc seconds, the OSA will call the ISA on call that night. That person will then:

  • Call the AO on-call person (to be specified beforehand by the AO group)
  • Negotiate with the observers the exact time for the hand over (e.g., after the current OB is complete.).
  • Notify team members in Germany via the luci email list
Special note regarding Test #1 = W-unit transform: This test must precede all other tests, and must be done around transit of the Stone M field at 18h 0d is up and observable. Note also, that, once collected, the data must be analyzed and implemented before proceeding with the other tests. Verification can be done with the same field or any of the other test fields

Planned tests:

  • W unit stage transform calibration (subsequent tests can be used to verify)
  • AO stability and high Strehl measurements (2 fields)
  • AFC stability (2 fields)
  • Strehl vs. radius (the AO or AFC stability data on globular clusters can be used for this)
Thanks in advance for your participation in this new commissioning model, and for your patience as we work out any bugs in these guidelines.

Log of interrupts (UTDate, Partner, Hours of interrupt, Results):
  • 20160530 LBTB 1.00hr Clouds rolled in ~15 minutes after handover - no data
  • 20160531 LBTB 2:00hr Problems getting AO system running - no data
  • 20160615 OSU/RC 2:00hr Problems with telescope focus - no data
  • 20160704 INAF 1:00hr Problems with binocular preset with DX Ad-Sec off-line - no data
  • 20160706 DDT 1:00hr AO and Preset problems - AO images for trouble-shooting

-- %USERSIG{DavidThompson - 2016-06-06}%

Comments


 
Topic revision: r6 - 06 Jul 2016, AlbertConrad
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback